Determine the original loan value loan amount for someone agreeing to make annual year end payments of

    $25,000/year for the next 20 years when the market rate is 4%. Note: you must use the pricing formula from class. Do not use 20 different terms.
Answer #1

Annual loan payment = 25,000

Payment Time period = 20 years

Interest Rate = 4%
Calculate the original loan value.
Here we have to calculate the Present value of the future annual payments.

PV = A (P/A, 4%, 20)

(P/A, I, N) factor is used to calculate the present value of annual uniform cash flows.
PV = 25,000(P/A, 4%, 20)

PV = 25000(13.59033) = 339,758.25

The original loan is $339,758.25

The formula used in calculating P/A factor is = (1+interest) ^n - 1 / interest (1+interest)^n